# NTK-DFL: Enhancing Decentralized Federated Learning in Heterogeneous Settings via Neural Tangent Kernel

### v1

@@ Abstract
Decentralized federated learning (DFL) is a collaborative machine learning framework for training a model across participants without a central server or raw data exchange. DFL faces challenges due to statistical heterogeneity, as participants often possess different data distributions reflecting local environments and user behaviors. Recent work has shown that the neural tangent kernel (NTK) approach, when applied to federated learning in a centralized framework, can lead to improved performance. The NTK-based update mechanism is more expressive than typical gradient descent methods, enabling more efficient convergence and better handling of data heterogeneity. We propose an approach leveraging the NTK to train client models in the decentralized setting, while introducing a synergy between NTK-based evolution and model averaging. This synergy exploits inter-model variance and improves both accuracy and convergence in heterogeneous settings. Our model averaging technique significantly enhances performance, boosting accuracy by at least 10% compared to the mean local model accuracy. Empirical results demonstrate that our approach consistently achieves higher accuracy than baselines in highly heterogeneous settings, where other approaches often underperform. Additionally, it reaches target performance in 4.6 times fewer communication rounds. We validate our approach across multiple datasets, network topologies, and heterogeneity settings to ensure robustness and generalizability.
@@ 1 Introduction
Federated learning (FL) is a machine learning paradigm in which multiple clients train a global model without the explicit communication of training data. In most FL scenarios, clients communicate with a central server that performs model aggregation. In the popular federated averaging (FedAvg) algorithm (McMahan et al., 2017), clients perform multiple rounds of stochastic gradient descent (SGD) on their own local data, then send this new weight vector to a central server for aggregation. As FL gains popularity in both theoretical studies and real-world applications, numerous improvements have been made to address challenges, including communication efficiency, heterogeneous data distributions, and security concerns (Sattler et al., 2020; Li et al., 2020; Zhu et al., 2019). To handle the performance degradation caused by data heterogeneity, many works have proposed mitigation for FedAvg (Karimireddy et al., 2020; Li et al., 2020). Notably, some researchers have introduced the neural tangent kernel (NTK), replacing the commonly-used SGD in order to improve the model convergence (Yu et al., 2022; Yue et al., 2022).
Despite these advancements, the centralized nature of traditional FL schemes introduces the possibility for client data leakage, computational bottlenecks at the server, and high communication bandwidth demand (Kairouz et al., 2021). Decentralized federated learning (DFL) has been proposed as a solution to these issues (Martínez Beltrán et al., 2023). In DFL, clients may communicate with each other along an undirected graph, where each node represents a client and each edge represents a communication channel between clients. While DFL addresses some of the issues inherent to centralized FL, both frameworks grapple with the challenge of statistical heterogeneity across clients. Although mixing data on a central server could readily resolve this issue, privacy concerns and the burden of extensive communication make FL and DFL approaches necessary to address this challenge. This paper focuses on the following research question: How can we design a DFL approach that effectively addresses statistical heterogeneity?
We propose a DFL method that exploits the NTK to evolve weights. We denote this paradigm NTK-DFL. Our approach combines the advantages of NTK-based optimization with the decentralized structure of DFL. The NTK-DFL weight evolution scheme makes use of the communication of client Jacobians, allowing for more expressive updates than traditional weight vector transmissions and improving performance under heterogeneity. Complementing this NTK-based evolution, we utilize a model averaging step that exploits inter-model variance among clients, creating a global model with much better generalization than any local model. The contributions of this paper are threefold.
@@
    The NTK-DFL aggregated model achieves at least 10% higher accuracy than the average accuracy of individual client models. This aggregated model exhibits robust performance across various network topologies, datasets, and data distributions.
@@ 2 Related Work
Decentralized Federated Learning (DFL) DFL aims to eliminate the need for a central server by connecting clients in a fully decentralized topology. Sun et al. (2023) adapted the FedAvg approach of multiple local SGD iterations to the decentralized setting, leveraging momentum to improve model convergence and weight quantization to reduce total communication cost. Our NTK-DFL method may be viewed as building on this foundation, using the neural tangent kernel for more effective weight updates. Dai et al. (2022) proposed a method of DFL where each client possesses their own sparse mask personalized to their specific data distribution. Shi et al. (2023) employed the sharpness-aware minimization optimizer to reduce the inconsistency of local models, whereas we tackle this issue through per-round averaging and final model aggregation. DFL approaches can aim to train one global model, such as the case of many hospitals training a model for tumor classification with local, confidential images (Shiri et al., 2022). They may also aim to train a personalized model for each client in order to perform better on the local data distribution. For example, different groups of mobile phone users may use different words or emojis and would benefit from a personalized model (Tan et al., 2023). Our method focuses on training a high-performing global model that generalizes well across all clients, offering improved convergence and resilience to data heterogeneity compared to existing DFL approaches.
Neural Tangent Kernel (NTK) NTK has primarily been used for the analysis of neural networks (Golikov et al., 2022), though it has recently seen use in the training of neural networks for FL (Yue et al., 2022). Introduced by Jacot et al. (2018), it shows that the evolution of an infinitely wide neural network converges to a kernelized model. This approach has enabled the analytical study of models that are well approximated by this infinite width limit (Liu et al., 2020). The NTK has also been extended to other model types, such as the recurrent neural network (Alemohammad et al., 2021) and convolutional neural network (Arora et al., 2019). We instead use the linearized model of the NTK approximation as a tool for weight evolution. Some studies have explored the integration of NTKs with FL. For instance, Huang et al. (2021) applied the NTK analysis framework to study the convergence properties of FedAvg, while Yu et al. (2022) extended NTK applications beyond theoretical analysis by training a convex neural network. Moreover, Yue et al. (2022) replaced traditional SGD-based optimization with NTK-based evolution in a federated setting, where clients transmit Jacobian matrices to a central server that performs weight updates using the NTK.
@@ 3.1 Problem Statement
We begin with a brief overview of centralized FL. The goal of centralized FL is to train a global model ${\bm{w}}$ across $M$ clients with their private, local data $\mathcal{D}_{i}=\{({\bm{x}}_{i,j},{\bm{y}}_{i,j})\}_{j=1}^{N_{i}}$, where $N_{i}$ is the number of training examples of the $i$th client. FL algorithms aim to numerically solve the sample-wise optimization problem of $\operatorname*{min}_{\bm{w}}F({\bm{w}})$, where $F({\bm{w}})=\frac{1}{M}\sum_{i=1}^{M}N_{i}F_{i}({\bm{w}})$ and $F_{i}({\bm{w}})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\mathcal{\ell}({\bm{w}},{\bm{x}}_{j},{\bm{y}}_{j})$, where $M$ denotes the number of clients. In the decentralized setting, an omnipresent global weight ${\bm{w}}$ is not available to clients in each communication round. Rather, each client possesses their own model ${\bm{w}}_{i}$ that is trained in the update process. Following related DFL work (Shi et al., 2023; Sun et al., 2023), we seek a global model ${\bm{w}}$ that benefits from the heterogeneous data stored locally across clients and generalizes better than any individual client model. A global or aggregated model may take the form ${\bm{w}}=\frac{1}{M}\sum_{i=1}^{M}N_{i}{\bm{w}}_{i}$.
@@
Notation Formally, we have a set of clients $\mathcal{C}=\{C_{1},C_{2},\ldots,{C_{M}}\}$. Each client is initialized with its weight ${\bm{w}}_{i}^{(0)}\in\mathbb{R}^{d}$, where $d$ is the size of the parameter vector and the superscript in ${\bm{w}}_{i}^{(0)}$ denotes the initial communication round. Model training is done in a series of communication rounds denoted $k\in\{1,2,...,K\}$. Let the graph at round $k$ be $\mathcal{G}^{(k)}=(\mathcal{C},E^{(k)})$, where $E^{(k)}$ is the set of edges representing connections between clients. Furthermore, the neighborhood of client $i$ at round $k$ is denoted $\mathcal{N}_{i}^{(k)}=\{C_{j}\mid(C_{i},C_{j})\in E^{(k)}\}$. This graph is specified before each communication round and can take an arbitrary form.
@@
Figure 1: NTK-DFL process:  Clients exchange weights,  Average weights with neighbors,  Compute and exchange Jacobians, labels, and function evaluations,  Construct local NTK and evolve weights [Eq. (5)]. This decentralized approach enables direct client collaboration and NTK-driven model evolution without a central server.
3.2 Communication Protocol
In the following sections, we present the NTK-DFL paradigm (Figure 1). We describe the key components of the algorithm, including the communication protocol and weight evolution process.
@@ 3.2 Communication Protocol
Per-round Parameter Averaging At the beginning of each communication round $k$, each client $C_{i}$ both sends and receives weights. Every client sends their model ${\bm{w}}_{i}^{(k)}$ to all neighbors $C_{j}\in\mathcal{N}_{i}^{(k)}$. Simultaneously, $C_{i}$ receives the weight vectors ${\bm{w}}_{j}^{(k)}$ from all neighbors $C_{j}\in\mathcal{N}_{i}^{(k)}$. Each client then aggregates its own weights along with the neighboring weights to form a new weight as follows:
$\bar{{\bm{w}}}_{i}^{(k)}=\frac{1}{|\mathcal{N}_{i}^{(k)}|+1}\big{(}{\bm{w}}_{i}^{(k)}+\sum_{j\in\mathcal{N}_{i}^{(k)}}{\bm{w}}_{j}^{(k)}\big{)}.$ (1)
@@
The client must then send this aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ back to all neighbors $C_{j}\in\mathcal{N}_{i}^{(k)}$. This step enables each client to construct a local NTK, comprised of inner products of Jacobians from both neighboring clients and their own Jacobians. See Algorithm 2 in Appendix A for a detailed description of this process.
@@
Local Jacobian Computation At this point, each client possesses its own aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ as well as an aggregated weight $\bar{{\bm{w}}}_{j}^{(k)}$ for each of their neighbors $C_{j}\in\mathcal{N}_{i}^{(k)}$. The clients use these weights $\bar{{\bm{w}}}_{j}^{(k)}$ and local data $\mathbf{X}_{i}$ to compute the Jacobian of $f(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$ with respect to the neighboring model parameters $\bar{{\bm{w}}_{j}}^{(k)}$ and their local data $\mathbf{X}_{i}$. We can denote this neighbor-specific Jacobian as
${\bm{J}}^{(k)}_{i,j}\equiv[\nabla_{\bar{{\bm{w}}_{j}}}{\bm{f}}(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})]^{\top}.$ (2)
@@
For a given client, the gradient is taken with respect to its neighbor’s aggregated weight $\bar{{\bm{w}}}_{j}^{(k)}$, but the function is evaluated on the client’s local data $\mathbf{X}_{i}$. Each client sends every neighbor their respective Jacobian tensor ${\bm{J}}^{(k)}_{i,j}$, true label $\mathbf{Y}_{i}$, and function evaluation ${\bm{f}}(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$. Note the order of the indices in the Jacobian: the client sends ${\bm{J}}_{i,j}^{(k)}$, an evaluation on the client’s data and the neighbor’s weights. In contrast, the client receives ${\bm{J}}_{j,i}^{(k)}$ from each of its neighbors, an evaluation on the neighbor’s data and the client’s weights. Algorithm 3 in Appendix A describes this process.
3.3 Weight Evolution
@@
After all inter-client communication is completed, the clients begin the weight evolution phase of the round (see Algorithm 4 in Appendix A). Here, all clients act in parallel as computational nodes. Each client possesses their own Jacobian tensor ${\bm{J}}_{i,i}^{(k)}$ as well as their neighboring Jacobian tensors ${\bm{J}}_{j,i}^{(k)}$ for each $C_{j}\in\mathcal{N}_{i}^{(k)}$.
@@
We denote the tensor of all Jacobian matrices possessed by a client at round $k$ as $\bm{\mathcal{J}}_{i}^{(k)}$, which is composed of matrices from the set $\{{\bm{J}}_{i,i}^{(k)}\}\cup\{{\bm{J}}_{j,i}^{(k)}\ |\ C_{j}\in\mathcal{N}_{i}^{(k)}\}$ stacked along a third dimension. We denote the matrix of true labels and function evaluations stacked in the same manner as $\bm{\mathcal{Y}}_{i}$ and ${\bm{f}}(\bm{\mathcal{X}}_{i})$, respectively. Here, $i$ denotes a client index $C_{i}\in\mathcal{C}$. Each tensor $\bm{\mathcal{J}}_{i}$, $\bm{\mathcal{Y}}_{i}$, and ${\bm{f}}(\bm{\mathcal{X}}_{i})$ is a stacked representation of the data from each client and its neighbors. Explicitly, we have $\bm{\mathcal{J}}_{i}^{(k)}\in\mathbb{R}^{\tilde{N_{i}}\times d_{2}\times d}$, $\bm{\mathcal{Y}}_{i}^{(k)}\in\mathbb{R}^{\tilde{N_{i}}\times d_{2}}$, and ${\bm{f}}(\bm{\mathcal{X}}_{i})\in\mathbb{R}^{\tilde{N_{i}}\times d_{2}}$. $\tilde{N_{i}}$ denotes the total number of data points between client $C_{i}$ and its neighbors, and $d_{2}$ is the output dimension.
@@
From here, each client performs the following operations to evolve its weights. First, compute the local NTK $\mathbf{H}_{i}^{(k)}$ from the Jacobian tensor $\bm{\mathcal{J}}_{i}^{(k)}$ using the definition of the NTK
$\mathbf{H}^{(k)}_{i,mn}=\frac{1}{d_{2}}\langle\bm{\mathcal{J}}_{i}^{(k)}({\bm{x}}_{m}),\bm{\mathcal{J}}_{i}^{(k)}({\bm{x}}_{n})\rangle_{F}.$ (3)
@@
Each element of the NTK is a pairwise Frobenius inner product between Jacobian matrices, where the indices $m$ and $n$ correspond to two different data points. Second, using $\mathbf{H}_{i}^{(k)}$, the client evolves their weights as follows:
@@
${\bm{f}}^{(k,t)}(\bm{\mathcal{X}}_{i})=(\mathbf{I}-e^{\frac{\eta t}{\tilde{N_{i}}}\mathbf{H}_{i}^{(k)}})\bm{\mathcal{Y}}_{i}^{(k)}+e^{\frac{\eta t}{\tilde{N_{i}}}\mathbf{H}_{i}^{(k)}}{\bm{f}}^{(k)}(\bm{\mathcal{X}}_{i}),$ (4)
${\bm{w}}_{i}^{(k,t)}=\sum_{j=1}^{d_{2}}(\bm{\mathcal{J}}^{(k)}_{i,:j:})^{\top}\mathbf{R}_{i,:j}^{(k,t)}+\bar{{\bm{w}}}_{i}^{(k)},\qquad\mathbf{R}_{i,:j}^{(k,t)}\equiv\frac{\eta}{\tilde{N_{i}}d_{2}}\sum_{u=0}^{t-1}[\bm{\mathcal{Y}}_{i}^{(k)}-{\bm{f}}^{(k,u)}(\bm{\mathcal{X}}_{i})].$ (5)
@@
Final Model Averaging  Throughout the paper, we study the convergence of the aggregated model ${\bm{w}}=\frac{1}{{M}}\sum_{i=1}^{M}N_{i}{\bm{w}}_{i}$. In the decentralized setting, clients would average all models to create ${\bm{w}}$ after all training is completed. This may be done through a fully-connected topology, sequential averaging on a ring topology, or in a secure, centralized manner. Clients may also connect in a denser topology than that of training, and average with a desired number of neighbors. In practice, we observe that the aggregated model is more accurate than any individual client model. We study the impact of client averaging order on model performance with a client selection algorithm and show the results in Figure 6. Each client that opts in to model averaging contributes a portion of its data to a global validation set before training begins. Our client selection algorithm selects clients in the order of their accuracy on the validation set. We will demonstrate that in the practical setting, with a proper selection of clients, not all nodes must opt into final model averaging in order for the aggregate model to benefit from improved convergence. We note a difference between model consensus, often discussed in the DFL literature (Savazzi et al., 2020; Liu et al., 2022), and the proposed final model averaging approach. Model consensus refers to the gradual convergence of all client models to a single, unified model over numerous communication rounds. In contrast, our approach implements final model averaging as a distinct step performed after the completion of the training process.

### v2

@@ Abstract
Decentralized federated learning (DFL) is a collaborative machine learning framework for training a model across participants without a central server or raw data exchange. DFL faces challenges due to statistical heterogeneity, as participants often possess data of different distributions reflecting local environments and user behaviors. Recent work has shown that the neural tangent kernel (NTK) approach, when applied to federated learning in a centralized framework, can lead to improved performance. We propose an approach leveraging the NTK to train client models in the decentralized setting, while introducing a synergy between NTK-based evolution and model averaging. This synergy exploits inter-client model deviation and improves both accuracy and convergence in heterogeneous settings. Empirical results demonstrate that our approach consistently achieves higher accuracy than baselines in highly heterogeneous settings, where other approaches often underperform. Additionally, it reaches target performance in 4.6 times fewer communication rounds. We validate our approach across multiple datasets, network topologies, and heterogeneity settings to ensure robustness and generalization. Source code for NTK-DFL is available at https://github.com/Gabe-Thomp/ntk-dfl.
Machine Learning, ICML, Federated Learning, Distributed Training, Decentralized Federated Learning, Neural Tangent Kernel
@@ 1 Introduction
Federated learning (FL) is a machine learning paradigm in which multiple clients train a global model without the explicit communication of training data. In most FL scenarios, clients communicate with a central server that performs model aggregation. In the popular federated averaging (FedAvg) algorithm (McMahan et al., 2017), clients perform multiple rounds of stochastic gradient descent (SGD) on their own local data, then send this new weight vector to a central server for aggregation. As FL gains popularity in both theoretical studies and real-world applications, numerous improvements have been made to address challenges, including communication efficiency, heterogeneous data distributions, and security concerns (Sattler et al., 2020; Li et al., 2020b; Zhu et al., 2019). To handle the performance degradation caused by data heterogeneity, many works have proposed mitigation strategies for FedAvg (Karimireddy et al., 2020; Li et al., 2020b). Notably, some researchers have introduced the neural tangent kernel (NTK), replacing the commonly-used SGD in order to improve the model convergence (Yu et al., 2022; Yue et al., 2022).
Despite these advancements, the centralized nature of traditional FL schemes introduces the possibility for client data leakage, computational bottlenecks at the server, and high communication bandwidth demand (Kairouz et al., 2021). Decentralized federated learning (DFL) has been proposed as a solution to these issues (Martínez Beltrán et al., 2023). In DFL, clients may communicate with each other along an undirected graph, where each node represents a client and each edge represents a communication channel between clients. While DFL addresses some of the issues inherent to centralized FL, both frameworks grapple with the challenge of statistical heterogeneity across clients. Although mixing data on a central server could readily resolve this issue, transmitting raw, private training data from clients introduces privacy concerns, making FL and DFL approaches good candidates to address this challenge (Yuan et al., 2024). This paper focuses on the following research question: How can we design a DFL approach that effectively addresses statistical heterogeneity?
We propose a method that exploits the NTK to evolve weights. We denote this paradigm NTK-DFL. Our approach combines the advantages of NTK-based optimization with the decentralized structure of DFL. The NTK-DFL weight evolution scheme makes use of the communication of client Jacobians, allowing for more expressive updates than traditional weight vector transmissions and improving performance under heterogeneity. Complementing this NTK-based evolution, we utilize a model averaging step that exploits inter-client model deviation, creating a global model with much better generalization than any local model. We demonstrate that NTK-DFL maintains high performance even under aggressive compression measures. Through reconstruction attack studies, we also analyze how this compression affects data privacy. The contributions of this paper are threefold.
@@
    The NTK-DFL aggregated model exhibits robust performance across various network topologies, datasets, data distributions, and compression measures. This performance is further supported by theoretical bounds demonstrating improved convergence rates.
@@ 2 Related Work
Decentralized Federated Learning (DFL) DFL aims to eliminate the need for a central server by connecting clients in a fully decentralized topology. Sun et al. (2021) adapted the FedAvg approach of multiple local SGD iterations to the decentralized setting. Our NTK-DFL method may be viewed as building on this foundation, using the neural tangent kernel for more effective weight updates. Dai et al. (2022) proposed a method of DFL where each client possesses their own sparse mask personalized to their specific data distribution. Shi et al. (2023) employed the sharpness-aware minimization optimizer to reduce the inconsistency of local models, whereas we tackle this issue through per-round averaging and final model aggregation. DFL approaches can aim to train one global model, such as the case of many hospitals training a model for tumor classification with local, confidential images (Shiri et al., 2022). They may also aim to train a personalized model for each client to perform better on the local data distribution. For example, different groups of smartphone users may use different emojis and would benefit from a personalized model (Tan et al., 2023). Our method focuses on training a high-performing global model that generalizes well across all clients, offering improved convergence and resilience to data heterogeneity compared to existing DFL methods.
Neural Tangent Kernel (NTK) NTK has primarily been used for the analysis of neural networks (Golikov et al., 2022), though it has recently seen use in the training of neural networks for FL (Yue et al., 2022). Introduced by Jacot et al. (2018), it shows that the evolution of an infinitely wide neural network converges to a kernelized model. This approach has enabled the analytical study of models that are well approximated by this infinite width limit (Liu et al., 2020). NTK has also been extended to other model types, such as the recurrent neural network (Alemohammad et al., 2021) and convolutional neural network (Arora et al., 2019). We instead use the linearized model of the NTK approximation as a tool for weight evolution. Some studies have explored the integration of NTKs with FL. For instance, Huang et al. (2021) applied the NTK analysis framework to study the convergence properties of FedAvg, while Yu et al. (2022) extended NTK applications beyond theoretical analysis by training a convex neural network. Moreover, Yue et al. (2022) replaced traditional SGD-based optimization with NTK-based evolution in a federated setting, where clients transmit Jacobian matrices to a central server that performs weight updates using NTK.
@@ 3.1 Problem Statement
We begin with a brief overview of centralized FL. The goal of centralized FL is to train a global model ${\bm{w}}$ across $M$ clients with their private, local data $\mathcal{D}_{i}=\{({\bm{x}}_{i,j},{\bm{y}}_{i,j})\}_{j=1}^{N_{i}}$, where $N_{i}$ is the number of training examples of the $i$th client. FL algorithms aim to numerically solve the sample-wise optimization problem of $\operatorname*{min}_{\bm{w}}F({\bm{w}})$, where $F({\bm{w}})=\frac{1}{M}\sum_{i=1}^{M}N_{i}F_{i}({\bm{w}})$ and $F_{i}({\bm{w}})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\mathcal{\ell}({\bm{w}};{\bm{x}}_{i,j},{\bm{y}}_{i,j})$.
In the decentralized setting, an omnipresent global weight ${\bm{w}}$ is not available to clients in each communication round. Rather, each client possesses their own model ${\bm{w}}_{i}$ that is trained in the update process. Following related DFL work (Shi et al., 2023; Sun et al., 2021), we seek a global model ${\bm{w}}$ that benefits from the heterogeneous data stored locally across clients and generalizes better than any individual client model ${\bm{w}}_{i}$. A global or aggregated model may take the form ${{\bm{w}}=\frac{1}{N}\sum_{i=1}^{M}N_{i}{\bm{w}}_{i}}$, where ${N=\sum_{i=1}^{M}N_{i}}$.
@@
Notation Formally, we have a set of clients $\mathcal{C}=\{1,\dots,i,\ldots,{M}\}$. Each client is initialized with its weight ${\bm{w}}_{i}^{(0)}\in\mathbb{R}^{d}$, where $d$ is the size of the parameter vector and the superscript in ${\bm{w}}_{i}^{(0)}$ denotes the initial communication round. Model training is done in a series of communication rounds denoted $k\in\{1,2,...,K\}$. Let the graph at round $k$ be $\mathcal{G}^{(k)}=(\mathcal{C},E^{(k)})$, where $E^{(k)}$ is the set of edges representing connections between clients. Furthermore, the neighborhood of client $i$ at round $k$ is denoted ${\mathcal{N}_{i}^{(k)}=\{j\mid(i,j)\in E^{(k)}\}}$. This graph is specified before each communication round and can take an arbitrary form.
@@
Figure 1: NTK-DFL process:  Clients exchange weights,  Average weights with neighbors,  Compute and exchange Jacobians, labels, and function evaluations,  Construct local NTK and evolve weights [Eq. (5a)]. This decentralized approach enables direct client collaboration and NTK-driven model evolution without a central server.
3.2 Proposed NTK-DFL
Figure 1 illustrates the proposed NTK-DFL method. We describe its key components below.
@@ 3.2 Proposed NTK-DFL
Per-round Parameter Averaging At the beginning of each communication round $k$, each client $i$ both sends and receives weights—client $i$ sends its model ${\bm{w}}_{i}^{(k)}$ to all neighbors ${j\in\mathcal{N}_{i}^{(k)}}$ and receives ${\bm{w}}_{j}^{(k)}$ from all neighbors. Each client then aggregates its own weights with its neighbors’ weights to form a new averaged weight as follows:
$\bar{{\bm{w}}}_{i}^{(k)}=\tfrac{1}{{N_{i}+\sum_{j\in\mathcal{N}^{(k)}_{i}}N_{j}}}\big{(}{N_{i}{\bm{w}}_{i}^{(k)}+\sum_{j\in\mathcal{N}_{i}^{(k)}}N_{j}{\bm{w}}_{j}^{(k)}}\big{)}.$ (1)
@@
The client must then send this aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ back to all neighbors. This step enables each client to construct a local NTK, comprised of inner products of Jacobians from both neighboring clients and its own Jacobian. See Algorithm 2 in Appendix A for details.
@@
Local Jacobian Computation At this point, each client possesses its own aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ as well as a set of aggregated weights $\bar{{\bm{w}}}_{j}^{(k)}$ from each of their neighbors ${j\in\mathcal{N}_{i}^{(k)}}$. The $i$th client computes the Jacobian of $f(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$ with respect to the neighboring model parameters $\bar{{\bm{w}}_{j}}^{(k)}$ using its local data $\mathbf{X}_{i}$. We denote this neighbor-specific Jacobian as
${\bm{J}}^{(k)}_{i,j}\equiv[\nabla_{{{\bm{w}}}}f(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})]^{\top}.$ (2)
@@
Each client sends every neighbor their respective Jacobian ${\bm{J}}^{(k)}_{i,j}$, true label $\mathbf{Y}_{i}$, and function evaluation $f(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$. Note the order of the indices in the Jacobian: the $i$th client sends ${\bm{J}}_{i,j}^{(k)}$, an evaluation on the $i$th client’s data and its neighbors’ weights. In contrast, the $i$th client receives ${\bm{J}}_{j,i}^{(k)}$ from each of its neighbors, an evaluation on its neighbors’ data and the $i$th client’s weights. Algorithm 3 in Appendix A describes this process.
@@
Weight Evolution After all inter-client communication is completed, the clients begin the weight evolution phase of the round (see Algorithm 4 in Appendix A). Here, all clients act in parallel as computational nodes. Each client possesses their own Jacobian tensor ${\bm{J}}_{i,i}^{(k)}$ as well as their neighboring Jacobian tensors ${\bm{J}}_{j,i}^{(k)}$ for each ${j\in\mathcal{N}_{i}^{(k)}}$.
@@
We denote the tensor of all Jacobian matrices possessed by the $i$th client at round $k$ as $\bm{\mathcal{J}}_{i}^{(k)}$, which is composed of matrices from the set $\{{\bm{J}}_{i,i}^{(k)}\}\cup\{{\bm{J}}_{j,i}^{(k)}\ |\ {j\in\mathcal{N}_{i}^{(k)}}\}$ stacked along the third dimension. We denote the matrix of true labels and function evaluations stacked in the same manner as $\bm{\mathcal{Y}}_{i}$ and ${\bm{f}}(\bm{\mathcal{X}}_{i})$, respectively. Explicitly, we have $\bm{\mathcal{J}}_{i}^{(k)}\in\mathbb{R}^{\tilde{N_{i}}\times d_{2}\times d}$, $\bm{\mathcal{Y}}_{i}^{(k)}\in\mathbb{R}^{\tilde{N_{i}}\times d_{2}}$, and ${\bm{f}}(\bm{\mathcal{X}}_{i})\in\mathbb{R}^{\tilde{N_{i}}\times d_{2}}$. Additionally, $\tilde{N_{i}}=N_{i}+\sum_{j\in\mathcal{N}_{i}^{(k)}}N_{j}$ represents the total number of data points between client $i$ and its neighbors, and $d_{2}$ is the output dimension.
@@
From here, each client performs the following operations to evolve its weights. First, compute the local NTK $\mathbf{H}_{i}^{(k)}$ from the Jacobian tensor $\bm{\mathcal{J}}_{i}^{(k)}$ using the definition of NTK:
$[\mathbf{H}^{(k)}_{i}]_{m,n}=\frac{1}{d_{2}}\langle\bm{\mathcal{J}}_{i}^{(k)}({\bm{x}}_{m}),\bm{\mathcal{J}}_{i}^{(k)}({\bm{x}}_{n})\rangle_{\text{F}}.$ (3)
@@
Each element of the NTK is a pairwise Frobenius inner product between Jacobian matrices, where the indices $m$ and $n$ correspond to the $m$th and $n$th data points, respectively. Second, using $\mathbf{H}_{i}^{(k)}$, the client evolves their weights as follows (see Appendix B for more details):
@@
$\begin{split}{\bm{f}}^{(k,t)}(\bm{\mathcal{X}}_{i})&=\bigl{(}\mathbf{I}\!-\!e^{-\frac{\eta t}{\tilde{N}_{i}}\mathbf{H}_{i}^{(k)}}\bigr{)}\bm{\mathcal{Y}}_{i}^{(k)}\\ &\quad+\,e^{-\frac{\eta t}{\tilde{N}_{i}}\mathbf{H}_{i}^{(k)}}{\bm{f}}^{(k)}(\bm{\mathcal{X}}_{i}).\end{split}$ (4)
We unroll gradient steps to find the weight ${\bm{w}}_{i}^{(k,t)}$ as follows:
${\bm{w}}_{i}^{(k,t)}$ $=\sum_{j=1}^{d_{2}}(\bm{\mathcal{J}}^{(k)}_{i,:j:})^{\top}\mathbf{R}_{i,:j}^{(k,t)}+\bar{{\bm{w}}}_{i}^{(k)},$ (5a)
$\mathbf{R}_{i,:j}^{(k,t)}$ $\equiv\frac{\eta}{\tilde{N_{i}}d_{2}}\sum_{u=0}^{t-1}[\bm{\mathcal{Y}}_{i}^{(k)}-{\bm{f}}^{(k,u)}(\bm{\mathcal{X}}_{i})].$ (5b)
@@
Final Model Averaging  Throughout the paper, we study the convergence of the final averaged model $\bar{{\bm{w}}}^{(k)}=\frac{1}{M}\sum_{i=1}^{M}{\bm{w}}_{i}^{(k)}$. In the decentralized setting, clients would average all $K$ client models after all training is completed. This may be done through a fully-connected topology, sequential averaging on a ring topology, or in a secure, centralized manner. The average may also be carried out over a subset of clients. In practice, we observe that the aggregated model $\bar{{\bm{w}}}^{(k)}$ generally performs better than any individual client model ${{\bm{w}}}_{i}^{(k)}$. We study the impact of the order of client averaging on model performance with a client selection algorithm and show the results in Figure 7. Each client that opts in to model averaging contributes a portion of its data to a global validation set before training begins. Our client selection algorithm selects clients in the order of their accuracy on the validation set. We will demonstrate that in the practical setting, with a proper selection of clients, not all nodes must opt into final model averaging in order for the aggregated model to benefit from improved convergence. We note a difference between model consensus, often discussed in the DFL literature (Savazzi et al., 2020; Liu et al., 2022), and the proposed final model averaging approach. Model consensus refers to the eventual convergence of all client models to a single, unified model over numerous communication rounds. In contrast, our approach implements final model averaging as a distinct step performed after the completion of the training process.